OpenStudy (anonymous):

Solve. negative two-fifths times x – 9 < nine-tenths

OpenStudy (karatechopper):

\[-\frac{ 2 }{ 5 } \times x-9 < \frac{ 9 }{ 10 }\] U want me to solve that?